core: 0 of 1 pt ocial Sci 14.5.21 Sociologists have found that crime rates are infiuenced by tomperature. In a town of 200,000 peoplo, the crime rato has been approximtd 5 of 8 (2 complete) ? Score: 20%, 2 of 10 pts Question Help T-65 +100, where C is the number of crimes per month and T is the average monthly temperature in degrees Fahrenheit. The average temperature for May was 75 and by the end of May the temperature was rising at the rate of 6 per month. How fast is the crime rate rising at the end of May? Atthe end of May, the crime rate is rising by?crime(s) per month. Simplify your answer.)
Answer #1

In a town of 200000 people, the crime rate ( C ) is given by C = (1/10)(T-65)2 +100, where T is the average temperature in 0 F during the month. Further, the average temperature during the month of May was 750 F so that the crime rate in May was (1/10) (75-65)2+100 = (1/10)*100+ 100 = 110. Also, the average temperature during the month of June was (75+6)0 F=810 F so that the crime rate in June was (1/10)(81-65)2 +100 =(1/10)*256 +100 = 125.6. Thus, the rise in crime rate at the end of May is [(125.6-110)/110]* 100 % = 14.18 % ( on rounding off to 2 decimal places).